# PlainQAFact: Retrieval-augmented Factual Consistency Evaluation Metric for Biomedical Plain Language Summarization

### v2

@@ Abstract
Hallucinated outputs from large language models (LLMs) pose risks in the medical domain, especially for lay audiences making health-related decisions. Existing automatic factual consistency evaluation methods, such as entailment- and question-answering (QA) -based, struggle with plain language summarization (PLS) due to elaborative explanation phenomenon, which introduces external content (e.g., definitions, background, examples) absent from the scientific abstract to enhance comprehension. To address this, we introduce PlainQAFact, an automatic factual consistency evaluation metric trained on a fine-grained, human-annotated dataset PlainFact, for evaluating factual consistency of both source-simplified and elaborately explained sentences. PlainQAFact first classifies sentence type, then applies a retrieval-augmented QA scoring method. Empirical results show that existing evaluation metrics fail to evaluate the factual consistency in PLS, especially for elaborative explanations, whereas PlainQAFact consistently outperforms them across all evaluation settings. We further analyze PlainQAFact’s effectiveness across external knowledge sources, answer extraction strategies, answer overlap measures, and document granularity levels, refining its overall factual consistency assessment. Taken together, our work presents the first evaluation metric designed for PLS factual consistency evaluation, providing the community with both a robust benchmark and a practical tool to advance reliable and safe plain language communication in the medical domain. PlainQAFact and PlainFact are available at: https://github.com/zhiwenyou103/PlainQAFact.
@@ 1 Introduction
To address these challenges, we propose a targeted retrieval-based metric for factual consistency evaluation in PLS (Figure 1). We introduce a new expert-annotated dataset, PlainFact, consisting of human-written plain language summaries aligned with scientific abstracts. Each sentence is labeled with its type (elaborative explanation vs. source simplification), functional role, and alignment to the source (§3.1). Building on this, we present PlainQAFact, a dual-stage QA-based evaluation metric that selectively applies retrieval only to elaborative explanations (§3.2). This design ensures both efficiency and fidelity in evaluating factual consistency. Experiment results on several PLS datasets demonstrate the effectiveness of our approach, particularly in assessing complex, elaborative content (4, §5). Taken together, our work presents the first evaluation metric designed for PLS factual consistency evaluation, providing the community with both a robust benchmark and a practical tool to advance reliable and safe plain language communication in the medical domain.
@@ 3.1.2 Expert Annotation
Since each summary-abstract pair is authored by the same individual, we assume all information to be factual. The annotation aims to capture how plain language sentences relate to their scientific abstracts. Annotators analyze each plain language sentence across three dimensions: (1) Factuality type: identify whether a sentence is a source simplification (derived from the abstract) or an elaborative explanation (introducing new content); (2) Functional role: categorize the sentence as background, definition, example, method/result, or other; and (3) Sentence alignment: map each plain language sentence to its corresponding sentence(s) in the scientific abstract. Details of annotation guidelines are provided in A.
Annotations are conducted by four independent annotators, each with at least a bachelor’s degree in biomedical sciences and prior fact-checking experience. Annotators are recruited via Upwork and compensated from \$15 to \$20 per hour. Each summary-abstract pair is annotated by two independent annotators, with disagreements resolved by a third. Inter-rater agreement, measured by Cohen’s Kappa for factuality type, functional role, and sentence alignment are 0.43, 0.60, and 0.55 respectively, indicated moderate agreement for all tasks [33].
Table 1 summarizes the dataset characteristics. Notably, 44% of plain language sentences are elaborative explanations, highlighting their role in enhancing the readability of plain language summaries. 66% of these cannot be directly verified against the source abstract, which underscores the need for factuality evaluation methods that account for such phenomena. Moreover, elaborative explanations include more background, definitions, and examples than source simplifications. For annotation examples, see B.
@@ 3.2.1 Learned Factuality Type Classifier
As elaborative explanations are prevalent in plain language generation and existing metrics struggle to capture added information [12], we first fine-tune a pre-trained language model to categorize factuality types of plain language sentences. Based on the factuality type annotations (source simplification vs. elaborative explanation) in PlainFact, we fine-tune the PubMedBERT-base model [^3] https://huggingface.co/microsoft/BiomedNLP-BiomedBERT-base-uncased-abstract-fulltext as a classifier. Our PlainFact is split 8:1:1 for training, validation, and testing. Additionally, we compare the pre-trained classifier with GPT-4o [34] as a zero-shot classifier (prompts are provided in C) to explore extending our factual consistency evaluation metric to domains and tasks lacking human-annotated data.
@@ 3.2.2 Domain Knowledge Retrieval
Retrieval-augmented methods have proven effective for explanation generation [28, 35], making them a natural fit for evaluating elaborative explanations. Since our dataset is in the medical domain, we employ MedCPT [36], a retriever pre-trained on large-scale PubMed search logs to generate biomedical text embeddings. For external resources, we incorporate StatPearls [37] for clinical decision support and medical textbooks [38] for domain-specific knowledge.
@@ 3.2.3 QA Evaluation Components
Gold Answer Extraction
The first step in QA-based factual consistency evaluation is to extract answer entities (keyphrases) from plain language summaries as gold answer, then verify factuality by comparing them with answers generated by a QA model for the same questions. If the generated answer is correct or relevant, the summary is considered factual. The previous study used PromptRank [40], a keyphrase generation method based on the T5 model. To compare answer extraction strategies, we use PromptRank as the baseline, and evaluate LLM-based extractors, including an open-source Llama 3.1 8B Instruct model (Llama 3.1) [41] and close-source GPT-4o (GPT-4o) [34]. Prompts for Llama 3.1 and GPT-4o are in C.
@@
We evaluate the alignment between the gold and generated answers from the QA model using the baseline Learned Evaluation metric for Reading Comprehension (LERC) score [45] and BERTScore [46]. The final step in PlainQAFact is weighting the factual consistency scores for source simplification and elaborative explanation sentences. Specifically, the simplification score $s$ is computed using only abstracts as the source in QA, while the explanation score $e$ is calculated by incorporating both abstracts and retrieved knowledge as source for QA-based evaluation.
@@ 4.1 Datasets
As introduced in Sec 3.1, we create PlainFact with fine-grained sentence-level annotations regarding the factuality types. Additionally, to test the generalizability of our pre-trained classifier, we also collect another subset from the test set of CELLS dataset [28], gathering 200 summary-abstract pairs. Similar to PlainFact3.1.1), we also conduct sentence-level perturbation for each plain language summary in our collected CELLS dataset. Therefore, we have 200 factual and 200 non-factual summary-abstract pairs for both PlainFact and CELLS datasets. We also compare PlainQAFact on the FareBio dataset, which contains 175 plain language summaries generated by seven different LLMs, together with the source articles [47]. Specifically, FareBio includes 25 full scientific article examples, each source article is summarized by seven LLMs, resulting 175 plain language summaries in total. We sample plain language sentences that are not both faithful and factual as non-factual sentences, and the rest as factual sentences, resulting in 33 and 174 summaries respectively. We also extract sentences that are labeled as not faithful but factual in their “factual hallucination” label, which represents the plain language sentence that is not from the source article, but still conveying factual information, resulting in 53 valid summaries in total. These sentences are equal to our factual elaborative explanation sentences. Note in FareBio dataset, there are no non-factual simplification sentences. Therefore, we use the same set of non-factual sentences (33 sentences) for main experiment (Table 2) and explanation-only (Figure 2) evaluation.
@@ 4.2 Experiment Settings
The classifier used in PlainQAFact is fine-tuned on PlainFact with a batch size of 32 for 10 epochs. We apply early stopping during training based on validation loss. The temperature is set to 0 for GPT-4o and 0.01 for Llama 3.1. The Llama 3.1 model used in AE also uses a temperature of 0.01. For the QG model, we fine-tune BART-large with a batch size of 16, a learning rate of 3e-5, and two training epochs. The maximum input length is set to 512 for all models in PlainQAFact. More details of prompts and model hyper-parameters for GPT-4o and Llama 3.1 are provided in C and  D.
@@ 4.3 Existing Factuality Metrics
To the best of our knowledge, no prior work has efficiently evaluated factuality metrics for detecting elaborative explanations in PLS tasks. Moreover, most factuality metrics are designed for general-domain applications, largely due to limited quality and annotation of existing PLS datasets. To address this gap, we incorporate the following metrics in our experiment, informed by prior work [9]: (1) Dependency-Arc Entailment (DAE) [48] is an entailment-based method that evaluates summary factuality by breaking it into smaller entailment tasks at the arc level. The model independently determines whether each arc’s relationship is supported by the input. (2) AlignScore [19] is a model-based factuality metric using RoBERTa [49]. It extracts claims from the summary and calculates the alignment scores with all the context chunks from the input document. The final score is an average of all highest alignment probabilities. (3) SummaC [16] is an NLI-based inconsistency detection method designed for summarization tasks. It segments documents into sentences and aggregates scores between sentence pairs using a trained NLI model. (4) QAFactEval [39] is a QA-based factuality evaluation metric that assesses summary consistency by generating questions from the source document and comparing the model-generated answers with expected answers. (5) QuestEval [50] is a reference-less evaluation metric employs a T5 model to generate questions from the source document and verifies whether the summary can correctly answer them.
@@ 5 Results and Analysis
We first conduct a pilot study on the FactPICO dataset [9] to investigate the performance of existing factual consistency evaluation metrics on PLS tasks. Our study (G) reveals two major limitations. First, many LLM-generated plain language summaries in FactPICO contain incomplete or incoherent sentences, highlighting the need for a benchmark based on high-quality, human-authored plain language summaries. Second, existing automatic factual consistency metrics are not sensitive to evaluate the factuality of “added information”, underscoring the need for a more domain-related factual consistency evaluation metric, especially for those summaries with additional explanations that are not mentioned by the original source texts. To address these challenges, we introduce a new expert-annotated dataset, PlainFact, and propose PlainQAFact, a QA-based, retrieval-augmented metric for evaluating factual consistency in plain language summaries.
@@
Datasets Metrics Kendall’s $\tau$ Pearson AUC-ROC
CELLS Llama 3.1 57.5 42.1 83.4
GPT-4o 70.0 75.4 99.3
QAFactEval 61.6 70.6 93.5
QuestEval 23.2 28.3 66.4*
SummaC 24.8 29.8 67.5*
AlignScore 56.0 66.2 89.6
DAE 14.4 14.4 55.0*
PlainQAFact 62.1 75.2 93.8
FareBio Llama 3.1 40.5 38.9 83.0
GPT-4o 27.3 35.7 75.7
QAFactEval 32.8 43.9 81.6
QuestEval 37.8 45.2 86.4
SummaC 33.8 32.1 82.5
AlignScore 32.2 47.2 81.0
DAE 8.8 8.8 53.4
PlainQAFact 16.8 35.6 66.1
PlainFact Llama 3.1 60.7 52.6 85.3
GPT-4o 69.8 80.3 99.2
QAFactEval 65.8 79.3 96.4
QuestEval 28.4 34.7 70.1*
SummaC 34.1 42.8 74.1*
AlignScore 60.2 72.6 92.5
DAE 5.3 5.3 51.8*
PlainQAFact 65.7 81.3 96.4
Table 2: Evaluation results of automatic metrics on the CELLS [28], PlainFact, and FareBio [47] datasets. Results are evaluated using Kendall’s $\tau$, Pearson correlation, and AUC-ROC, with eight metric scores compared against human-labeled factuality. The standard deviations (std.) over five runs are: 0.2 (PlainQAFact), 0.5 (Llama 3.1), and 6.0 (GPT-4o) on PlainFact; 0.1 (PlainQAFact), 0.2 (Llama 3.1), and 6.3 (GPT-4o) on CELLS; and 0.1 (PlainQAFact), 2.9 (Llama 3.1), and 17.0 (GPT-4o) on FareBio. * indicates an improvement of our metric over prior work with 95% confidence interval (details in  E). Note the plain language summaries in FareBio are generated by seven LLMs, while CELLS and PlainFact are written by the authors of original articles.
@@ 5.1 Main Results
On CELLS, PlainQAFact outperforms all other automatic metrics on all three criteria (Kendall’s $\tau$=62.1, Pearson=75.2, AUC-ROC=93.8). This shows that our approach, first detecting elaborative content and then verifying it with retrieval, works better than existing QA-based metrics, which ask all questions from the source only, and also better than NLI-based evaluation metrics. Similarly, on PlainFact, PlainQAFact achieves the highest Pearson (81.3), surpassing GPT-4o by 1.0 point. It ties AUC-ROC (96.4) score with QAFactEval and has a competitive Kendall’s $\tau$ (65.7). These results on both datasets show that PlainQAFact is reliable for evaluating both factual and non-factual plain language summaries.
On FareBio, no single method performs the best on all criteria. QuestEval achieves the best AUC-ROC (86.4), AlignScore has the best Pearson (47.2), while Llama 3.1 reaches the best Kendall’s $\tau$ (40.5). However, FareBio differs from the other two datasets in several ways: (1) the original labels in FareBio are defined at the sentence level, and a summary is marked as non-factual if any sentence is non-factual; (2) their summaries are generated from full scientific articles rather than abstracts, making it harder for metrics to locate matching source text; (3) the dataset is imbalanced (33 non-factual vs. 174 factual); and (4) All plain language summaries are generated by LLMs and no elaborative explanation annotations are provided for each sentence or summary. These factors likely increase evaluation difficulty on FareBio and impact the performance of multiple metrics, including PlainQAFact and GPT-4o.
@@
Figure 2: Overall performance on human-annotated elaborative explanation summaries from PlainFact (392 summaries). The std. of PlainQAFact, Llama 3.1, and GPT-4o are 0.1, 1.0, and 7.7, respectively based on five runs for each metric. * indicates an improvement of our metric over prior work with 95% confidence interval (details in  E). PlainQAFact significantly outperforms most of the automatic factual consistency evaluation metrics in AUC-ROC. Note that the CELLS dataset does not contain annotations for elaborative explanations. Results of explanation-only evaluation on FactPICO and FareBio are reported in F.

### v3

@@ Abstract
Hallucinated outputs from large language models (LLMs) pose risks in the medical domain, especially for lay audiences making health-related decisions. Existing automatic factual consistency evaluation methods, such as entailment- and question-answering (QA) -based, struggle with plain language summarization (PLS) due to elaborative explanation phenomenon, which introduces external content (e.g., definitions, background, examples) absent from the scientific abstract to enhance comprehension. To address this, we introduce PlainQAFact, an automatic factual consistency evaluation metric trained on a fine-grained, human-annotated dataset PlainFact, for evaluating factual consistency of both source-simplified and elaborately explained sentences. PlainQAFact first classifies sentence type, then applies a retrieval-augmented QA scoring method. Empirical results show that existing evaluation metrics fail to evaluate the factual consistency in PLS, especially for elaborative explanations, whereas PlainQAFact consistently outperforms them across all evaluation settings. We further analyze PlainQAFact’s effectiveness across external knowledge sources, answer extraction strategies, answer overlap measures, and document granularity levels, refining its overall factual consistency assessment. Taken together, our work presents a sentence-aware, retrieval-augmented metric targeted at elaborative explanations in biomedical PLS tasks, providing the community with both a robust benchmark and a practical tool to advance reliable and safe plain language communication in the medical domain. PlainQAFact and PlainFact are available at: [Uncaptioned image] https://github.com/zhiwenyou103/PlainQAFact.
@@ 1 Introduction
To address these challenges, we propose a targeted retrieval-based metric for factual consistency evaluation in PLS (Figure 1). We introduce a new expert-annotated dataset, PlainFact, consisting of human-written plain language summaries aligned with scientific abstracts. Each sentence is labeled with its type (elaborative explanation vs. source simplification), functional role, and alignment to the source (§3.1). Building on this and prior studies for the QA-based evaluation metrics in the general domain [20, 21], we present PlainQAFact, a dual-stage QA-based evaluation metric that selectively applies retrieval only to elaborative explanations (§3.2). Experiment results on several PLS datasets demonstrate the effectiveness of our approach, particularly in assessing complex, elaborative content (§4, §5). Taken together, our work presents the first metric specifically designed to evaluate the factual consistency in elaborative explanations in PLS, providing the community with both a robust benchmark and a practical tool to advance reliable and safe plain language communication in the medical domain.
@@ 3.1.2 Expert Annotation
Since each summary-abstract pair is authored by the same individual, we assume all information to be factual. The annotation aims to capture how plain language sentences relate to their scientific abstracts. Annotators analyze each plain language sentence across three dimensions: (1) Factuality type: identify whether a sentence is a source simplification (derived from the abstract) or an elaborative explanation (introducing new content); (2) Functional role: categorize the sentence as background, definition, example, method/result, or other; and (3) Sentence alignment: map each plain language sentence to its corresponding sentence(s) in the scientific abstract. Details of annotation guidelines are provided in A. Note that “functional role” and “sentence alignment” are not components of developing PlainQAFact. We use these two fields in PlainFact as metadata to conduct error analysis and document our annotation process in A and  B.
Annotations are conducted by four independent annotators, each with at least a bachelor’s degree in biomedical sciences and prior fact-checking experience. Annotators are recruited via Upwork and compensated from \$15 to \$20 per hour. Each summary-abstract pair is annotated by two independent annotators, with disagreements resolved by a third. Inter-rater agreement (IRA), measured by Cohen’s Kappa for factuality type, functional role, and sentence alignment are 0.43, 0.60, and 0.55 respectively computed from 2,740 annotated plain language sentences, indicated fair agreement for all tasks [33].
Table 1 summarizes the dataset characteristics. Notably, 44% of plain language sentences are elaborative explanations, highlighting their role in enhancing the readability of plain language summaries. 66% of these cannot be directly verified against the source abstract, which underscores the need for factuality evaluation methods that account for such phenomena. Moreover, elaborative explanations include more background, definitions, and examples than source simplifications. For annotation examples, see B. Additionally, we provide qualitative analysis of PlainFact in  C given the fair IRA of factuality type.
@@ 3.2.1 Learned Factuality Type Classifier
As elaborative explanations are prevalent in plain language generation and existing metrics struggle to capture added information [12], we first fine-tune a pre-trained language model to categorize factuality types of plain language sentences. Based on the factuality type annotations (source simplification vs. elaborative explanation) in PlainFact, we fine-tune the PubMedBERT-base model [^3] https://huggingface.co/microsoft/BiomedNLP-BiomedBERT-base-uncased-abstract-fulltext as a classifier. Our PlainFact is split 8:1:1 for training, validation, and testing. Additionally, we compare the pre-trained classifier with GPT-4o [34] as a zero-shot classifier to explore extending our factual consistency evaluation metric to domains and tasks lacking human-annotated data.
@@ 3.2.2 Domain Knowledge Retrieval
Retrieval-augmented methods have proven effective for explanation generation [28, 35], making them a natural fit for evaluating elaborative explanations. Since our dataset is in the medical domain, we employ MedCPT [36], a retriever pre-trained on large-scale PubMed search logs to generate biomedical text embeddings. For external courpus, we incorporate StatPearls [^4] https://huggingface.co/datasets/MedRAG/statpearls [37] for clinical decision support and medical Textbooks [^5] https://huggingface.co/datasets/MedRAG/textbooks [38] for domain-specific medical knowledge (see details in  D.
@@ 3.2.3 QA Evaluation Components
Gold Answer Extraction (AE)
The first step in QA-based factual consistency evaluation is to extract answer entities (keyphrases) from plain language summaries as gold answer, then verify factuality by comparing them with answers generated by a QA model for the same questions. If the generated answer is correct or relevant, the summary is considered factual. The previous study used PromptRank [40], a keyphrase generation method based on the T5 model. To compare answer extraction strategies, we use PromptRank as the baseline, and evaluate LLM-based extractors, including an open-source Llama 3.1 8B Instruct model (Llama 3.1) [41] and close-source GPT-4o (GPT-4o) [34]. Prompts for Llama 3.1 and GPT-4o are in E.4.
@@
We evaluate the alignment between the gold and generated answers from the QA model using the baseline Learned Evaluation metric for Reading Comprehension (LERC) score [45] and BERTScore [46]. The final step in PlainQAFact is weighting the factual consistency scores for source simplification and elaborative explanation sentences. Specifically, the simplification score $s$ is computed using only abstracts as the source in QA, while the explanation score $e$ is calculated by incorporating both abstracts and retrieved knowledge as source for QA-based evaluation. $n$ is the number of sentences in a plain language summary, after we split the summary using Natural Language Toolkit (NLTK) [^7] https://www.nltk.org/ and and classify each sentence with our fine-tuned classifier.
@@ 4.1 Datasets
As introduced in Sec 3.1, we create PlainFact with fine-grained sentence-level annotations regarding the factuality types. Additionally, to test the generalizability of our pre-trained classifier, we also collect another subset from the test set of CELLS dataset [28], gathering 200 summary-abstract pairs. Similar to PlainFact3.1.1), we also conduct sentence-level perturbation for each plain language summary in our collected CELLS dataset. Therefore, we have 200 factual and 200 non-factual summary-abstract pairs for both PlainFact and CELLS datasets. We also compare PlainQAFact on the FareBio dataset, which contains 175 plain language summaries generated by seven different LLMs, together with the source articles [47]. Specifically, FareBio includes 25 full scientific articles, each summarized by seven LLMs, resulting 175 plain language summaries. Because FareBio does not include non-factual simplification sentences, we treat all simplification sentences as “factual.” For each original summary, we form new plain language summaries by aggregating all sentences annotated as factual and all sentences annotated as non factual separately. This produces 33 non factual and 174 factual summaries. We also identify explanation sentences that receive the “factual hallucination” label (i.e., they are not verifiable from the source but are factually correct based on external knowledge) and build 53 factual explanation only summaries from them. We use the same set of 33 non factual summaries for both the main experiment (Table 2) and the explanation only evaluation (Figure 6).
@@ 4.2 Experiment Settings
The classifier used in PlainQAFact is fine-tuned on PlainFact with a batch size of 32 for 10 epochs. We apply early stopping during training based on validation loss. The temperature is set to 0 for GPT-4o and 0.01 for Llama 3.1. The Llama 3.1 model used in AE also uses a temperature of 0.01. For the QG model, we fine-tune BART-large with a batch size of 16, a learning rate of 3e-5, and two training epochs. The maximum input length is set to 512 for all models in PlainQAFact. More details of prompts and model hyper-parameters for GPT-4o and Llama 3.1 are provided in E.
@@ 4.3 Existing Factuality Metrics
To the best of our knowledge, no prior work has efficiently evaluated factuality metrics for detecting elaborative explanations in PLS tasks. Moreover, most factuality metrics are designed for general-domain applications, largely due to limited quality and annotation of existing PLS datasets. To address this gap, we incorporate the following metrics in our experiment, informed by prior work [9]: (1) Dependency-Arc Entailment (DAE) [48] is an entailment-based method that evaluates summary factuality by breaking it into smaller entailment tasks at the arc level. The model independently determines whether each arc’s relationship is supported by the input. (2) AlignScore [19] is a model-based factuality metric using RoBERTa [49]. It extracts claims from the summary and calculates the alignment scores with all the context chunks from the input document. The final score is an average of all highest alignment probabilities. (3) SummaC [16] is an NLI-based inconsistency detection method designed for summarization tasks. It segments documents into sentences and aggregates scores between sentence pairs using a trained NLI model. (4) QAFactEval [39] is a QA-based factuality evaluation metric that assesses summary consistency by generating questions from the source document and comparing the model-generated answers with expected answers. (5) QuestEval [50] is a reference‐less evaluation metric employs a T5 model to generate questions from the source document and verifies whether the summary can correctly answer them.
@@ 5 Results and Analysis
We first conduct a pilot study on the FactPICO dataset [9] to investigate the performance of existing factual consistency evaluation metrics on PLS tasks. Our pilot study (F) reveals two major limitations. First, many LLM-generated plain language summaries in FactPICO contain incomplete or incoherent sentences, highlighting the need for a benchmark based on high-quality, human-authored plain language summaries. Second, existing automatic factual consistency metrics are not sensitive to evaluate the factuality of “added information,” underscoring the need for a more domain-related factual consistency evaluation metric, especially for those summaries with additional explanations that are not mentioned by the original source texts. To address these challenges, we introduce a new expert-annotated dataset, PlainFact, and propose PlainQAFact, a QA-based, retrieval-augmented metric for evaluating factual consistency in plain language summaries.
@@
Datasets Metrics Kendall’s $\tau$ Pearson AUC (99% CI)
CELLS Llama 3.1 57.5 42.1 83.4* [78.4, 88.0]
GPT-4o 70.0 75.4 99.3 [98.6, 99.8]
QAFactEval 61.6 70.6 93.5 [90.2, 96.3]
QuestEval 23.2 28.3 66.4* [59.1, 73.3]
SummaC 24.8 29.8 67.6* [60.3, 74.3]
AlignScore 56.0 66.2 89.6 [85.3, 93.3]
DAE 14.4 14.4 55.0* [50.7, 59.6]
PlainQAFact 62.1 75.2 93.8 [90.3, 96.7]
FareBio Llama 3.1 40.5 38.9 83.0 [69.0, 94.0]
GPT-4o 27.3 35.7 75.7 [59.7, 89.8]
QAFactEval 32.8 43.9 81.6 [69.2, 91.7]
QuestEval 37.8 45.2 86.4 [78.0, 93.3]
SummaC 33.8 32.1 82.5 [69.7, 92.9]
AlignScore 32.2 47.2 81.0 [65.9, 93.0]
DAE 8.8 8.8 53.4 [47.6, 57.6]
PlainQAFact 16.8 35.6 66.1 [48.8, 82.1]
PlainFact Llama 3.1 60.7 52.6 85.3* [80.6, 89.7]
GPT-4o 69.8 80.3 99.2 [98.4, 99.8]
QAFactEval 65.8 79.3 96.4 [94.0, 98.4]
QuestEval 28.4 34.7 70.1* [63.1, 76.5]
SummaC 34.1 42.8 74.1* [67.6, 80.4]
AlignScore 60.2 72.6 92.5 [93.8, 98.4]
DAE 5.3 5.3 51.7* [47.5, 55.9]
PlainQAFact 65.7 81.3 96.4 [93.6, 98.5]
Table 2: Evaluation results of automatic metrics on the CELLS [28], PlainFact, and FareBio [47] datasets. Results are evaluated using Kendall’s $\tau$, Pearson correlation, and AUC-ROC (AUC), with eight metric scores compared against human-labeled factual consistency. The standard deviations (std.) over five runs are: 0.2 (PlainQAFact), 0.5 (Llama 3.1), and 6.0 (GPT-4o) on PlainFact; 0.1 (PlainQAFact), 0.2 (Llama 3.1), and 6.3 (GPT-4o) on CELLS; and 0.1 (PlainQAFact), 2.9 (Llama 3.1), and 17.0 (GPT-4o) on FareBio. * indicates a statistically significant difference compared to PlainQAFact ($p<0.01$). Note the plain language summaries in FareBio are generated by seven LLMs, while CELLS and PlainFact are written by the authors of original articles.
@@ 5.1 Main Results
On CELLS, PlainQAFact outperforms all other automatic metrics on all three criteria. Specifically, PlainQAFact achieves the best Kendall’s $\tau$ ($62.1$) and Pearson ($75.2$) among all the automatic factual consistency evaluation metrics. As a comparative effect size, the pairwise difference in AUC-ROC between PlainQAFact and QAFactEval is small ($\Delta\mathrm{AUC-ROC}=0.3$, where $\Delta$ is PlainQAFact minus QAFactEval). After Holm–Bonferroni correction [52] on AUC-ROC, PlainQAFact shows significantly higher AUC-ROC than Llama 3.1, QuestEval, SummaC, DAE, and PromptRank ($p<0.01$), whereas the differences with QAFactEval, AlignScore, and GPT-4o are not statistically significant (adjusted $p\geq 0.91$). This shows that our approach, first detecting elaborative content and then verifying it with retrieval, performs better than most of the existing reference-based metrics, which ask all questions from the source only. Similarly, on PlainFact, PlainQAFact achieves the highest linear correlation with human ratings among automatic metrics (Pearson= $81.3$, 99% CI [76.3, 85.4]), surpassing GPT-4o. It ties AUC-ROC (96.4) score with QAFactEval and has a competitive Kendall’s $\tau$ (65.7). As a comparative effect size, the pairwise difference in AUC-ROC between PlainQAFact and QAFactEval is negligible ($\Delta\mathrm{AUC-ROC}=-0.1$). These results on both datasets show that PlainQAFact is reliable for evaluating the factual consistency of plain language summaries. See  G for more details of statistical tests.
On FareBio, no single method is best across all criteria. QuestEval achieves the best Kendall’s $\tau$ ($37.8$) and AUC-ROC ($86.4$), while AlignScore achieves the best Pearson correlation ($47.2$). However, FareBio differs from the other two datasets in several important ways: (1) the original sentence labels in FareBio are incomplete. Any sentence that can be traced back to the source is labeled as factual, so the dataset does not include sentences labeled as non-factual simplifications. This design leads to unreliable final summary-level labels. For example, even if a summary is labeled as factual and all of its sentences are labeled as factual, some sentences may still contain factual errors; (2) the processed dataset is highly imbalanced (33 non-factual vs. 174 factual summaries); (3) FareBio reports a fair Cohen’s Kappa of $0.48$ computed from 34 plain language sentences without adjudication, which indicates notable annotation quality issues in their curated dataset; and (4) all plain language summaries are generated by LLMs, and the plain language sentences are not guaranteed to be correct, since the annotation only checks whether the sentence is a source simplification without post-processed non-factual perturbation. These factors likely increase the difficulty of evaluation on FareBio and affect the performance of factual consistency evaluation metrics, including PlainQAFact and GPT-4o.
We also provide a detailed error analysis of FareBio that explains the main limitations of both PlainQAFact and the dataset (§5.5). In addition, because the non-factual plain language summaries of PlainFact and CELLS are generated by GPT-4o3.1.1), they may contain detectable artifacts that are not representative of real PLS errors. To further test the robustness of our proposed metric, we conduct additional experiments using the heuristic entity swap approach introduced in Guo et al. [12] to create non-factual plain language summaries. Specifically, following the same experiment setting as APPLS [12], we replace entities in each plain language sentence using the KBIN method [53], which replaces specific entity spans (such as names of viruses or diseases) with related but different concepts derived from the Unified Medical Language System (UMLS) [^8] https://www.nlm.nih.gov/research/umls/ while maximizing NLI contradiction and minimizing LM perplexity. This approach ensures that each new perturbed sentence is not only different from the original sentence, but also logically contradicts it while remaining grammatically fluent. Overall, on the CELLS dataset, for both the LLM-perturbed setting and the heuristic entity swap method, our metric achieves the best performance across all three evaluation criteria compared with other automatic metrics, showing the robustness of our proposed metric. Detailed experiment results are reported in  H.
@@
Figure 2: Overall performance on human-annotated elaborative explanation summaries from PlainFact (392 summaries). The std. of PlainQAFact, Llama 3.1, and GPT-4o are 0.1, 1.0, and 7.7, respectively based on five runs for each metric. * indicates a statistically significant difference compared to PlainQAFact ($p<0.01$). PlainQAFact significantly outperforms most of the automatic factual consistency evaluation metrics in AUC-ROC. Note that the CELLS dataset does not contain annotations for elaborative explanations. Results of explanation-only evaluation on FactPICO and FareBio are reported in I.